OpenStudy (anonymous):

Which value of x makes the following equation true? 4.6(x - 4) = -0.4x + 6.6 5 4 8 7

OpenStudy (paki):

just put the given values in equation and see.... if it gives both sides equal, then it is true... otherwise not,....

OpenStudy (anonymous):

4.6(x -4) = -0.4x + 6.6 4.6x - 18.4 = -0.4x + 6.6 4.6x + 0.4x = 6.6 + 18.4 5x = 25 I suppose you'll know the answer from here :)